# Morita equivalence of Nijenhuis structures

### v1

@@ 2.3 Morita equivalence of Nijenhuis groupoids
is a cartesian (pullback) diagram, i.e., $H_{1}\cong(\varphi\times\varphi)^{*}G_{1}$. Two Lie groupoids $\mathscr{G}$ and $\mathscr{H}$ are Morita equivalent if and only if they are connected by a span of Morita morphisms $\mathscr{G}\leftarrow\mathscr{P}\to\mathscr{H}$. The diagram $\mathscr{G}\leftarrow\mathscr{P}\to\mathscr{H}$ is also refered as a Morita equivalence (in terms of Morita morphisms) between $\mathscr{G}$ and $\mathscr{H}$.
@@
    $\left(\mathscr{G},N_{1}\right)\xleftarrow{\quad\varphi_{1}\quad}\left(\mathscr{P},K\right)\xrightarrow{\quad\varphi_{2}\quad}\left(\mathscr{H},N_{2}\right),$ (3)
@@
    $\mathrm{d}\varphi_{1}\circ K=N_{1}\circ\mathrm{d}\varphi_{1}\qquad\text{and}\qquad\mathrm{d}\varphi_{2}\circ K=N_{2}\circ\mathrm{d}\varphi_{2}.$ (4)
@@
    $\left(G_{1},N_{1}\right)$$\left(P,J\right)$$\left(H_{1},N_{2}\right)$$\left(G_{0},r_{1}\right)$$\left(H_{0},r_{2}\right)$$\scriptstyle{\mu_{1}}$$\scriptstyle{\mu_{2}}$ (5)
@@
    $N_{1}\times J\times N_{2}\left(TG_{1}\times_{TG_{0}}TP\times_{TH_{0}}TH_{1}\right)\subseteq TG_{1}\times_{TG_{0}}TP\times_{TH_{0}}TH_{1}$ (6)
@@
    $J\circ\mathrm{d}\mathscr{A}=\mathrm{d}\mathscr{A}\circ\left(N_{1}\times J\times N_{2}\right),$ (7)
@@
$g\cdot\left(g^{\prime},p\right)=\left(gg^{\prime},p\right)\qquad\text{and}\qquad\left(g,p^{\prime}\right)\cdot p=\left(g\,\varphi_{1}(p),\mathbf{s}_{\mathscr{P}}(p)\right).$ (8)
@@
$\mathrm{d}\mathscr{A}_{1}\circ N_{1}\times(N_{1}\times r_{K})=\left(N_{1}\times r_{K}\right)\circ\mathrm{d}\mathscr{A}_{1}\qquad\text{and}$ (9)
$\mathrm{d}\widetilde{\mathscr{A}}_{1}\circ\left(N_{1}\times r_{K}\right)\times K=\left(N_{1}\times r_{K}\right)\circ\mathrm{d}\widetilde{\mathscr{A}}_{1}.$ (10)
@@
$p\cdot(p^{\prime},h)=\left(\mathbf{t}_{\mathscr{P}}(p),\varphi_{2}(p)\,h\right)\qquad\text{and}\qquad(p,h^{\prime})\cdot h=(p,h^{\prime}h).$ (11)
@@
$\mathrm{d}\widetilde{\mathscr{A}}_{2}\circ\left(K\times\left(r_{K}\times N_{2}\right)\right)=\left(r_{K}\times N_{2}\right)\circ\mathrm{d}\widetilde{\mathscr{A}}_{2}\qquad\text{and}$ (12)
$\mathrm{d}\mathscr{A}_{2}\circ\left(\left(r_{K}\times N_{2}\right)\times N_{2}\right)=\left(r_{K}\times N_{2}\right)\circ\mathrm{d}\mathscr{A}_{2}.$ (13)
@@
Conversely, assume there exists a principal bibundle (5) satisfying (6) and (7). Let $\mathscr{P}=\left(P_{1}\rightrightarrows P\right)$ be the Lie groupoid where $P_{1}\coloneqq G_{1}\times_{(\mathbf{s}_{\mathscr{G}},\mu_{1})}P\times_{(\mu_{2},\mathbf{t}_{\mathscr{H}})}H_{1}$. The source and target maps are $(g,p,h)\mapsto p$ and $(g,p,h)\mapsto g\cdot p\cdot h$, respectively, while the multiplication is $((g,p,h),(g^{\prime},p^{\prime},h^{\prime}))\mapsto(gg^{\prime},p^{\prime},hh^{\prime})$. Then the diagram
@@
$\mathrm{d}\mu_{1}\circ J=r_{1}\circ\mathrm{d}\mu_{1}\qquad\text{and}\qquad\mathrm{d}\mu_{2}\circ J=r_{2}\circ\mathrm{d}\mu_{2}.$ (14)
@@
$N_{1}\times J\left(TG_{1}\times_{TG_{0}}TP\right)\subseteq TG_{1}\times_{TG_{0}}TP,\qquad J\times N_{2}\left(TP\times_{TH_{0}}TH_{1}\right)\subseteq TP\times_{TH_{0}}TH_{1},$ (15)
@@
$J\circ\mathrm{d}\mathscr{A}_{1}=\mathrm{d}\mathscr{A}_{1}\circ\left(N_{1}\times J\right),\qquad J\circ\mathrm{d}\mathscr{A}_{2}=\mathrm{d}\mathscr{A}_{2}\circ\left(J\times N_{2}\right).$ (16)
@@
2.4 Morita equivalence is an equivalence relation.
@@
$P\diamond Q\coloneqq\frac{P\times_{H_{0}}Q}{\mathscr{H}},$ (17)
@@
For reflexivity, let $\mathscr{G}$ be a Lie groupoid equipped with a multiplicative $(1,1)$-tensor field $N\in\Omega^{1}(\mathscr{G},T\mathscr{G})$. Because the actions of the principal bibundle $G_{0}\xleftarrow{\ \mathbf{t}\ }G_{1}\xrightarrow{\ \mathbf{s}\ }G_{0}$ correspond to left and right multiplication, and because $N$ is multiplicative, the tensor fields $N$ and $N\times N\times N$ satisfy (7).
@@
3 Infinitesimal Nijenhuis structures
In this section, we study the infinitesimal version of the notion of Morita equivalence for Nijenhuis groupoids that we developed in Section 2. We begin by defining infinitesimal Nijenhuis structures and providing an equivalent characterization in terms of $1$-derivations. We define the notion of Morita equivalence of infinitesimal Nijenhuis structures in Definition 3.12, and in Theorem 3.18 we show that it corresponds to the differentiation of the global Morita equivalence of Nijenhuis groupoids.
@@ 3 Infinitesimal Nijenhuis structures
An infinitesimal Nijenhuis structure is a pair $(A,R)$, where $A\to M$ is a Lie algebroid and $R\in\Omega^{1}(A,TA)$ is a Nijenhuis tensor field on the total space of $A$, such that $R:TA\to TA$ is a morphism of Lie algebroids from the tangent Lie algebroid to itself.
@@
for some $r\in\Omega^{1}(M,TM)$. Linear $(1,1)$-tensor fields that are morphisms of Lie algebroids are the infinitesimal counterparts of multiplicative $(1,1)$-tensors, they are called IM (infinitesimally multiplicative) $(1,1)$-tensor fields [9]. These objects can be characterized in algebraic terms as $1$-derivations as we explain in the following subsection.
3.1 Characterization of IM tensor fields via 1-derivations
A $1$-derivation on a vector bundle $\pi:E\to M$ is a triple $\mathscr{D}=\left(\nabla,\ell,r\right)$, where $r\in\Omega^{1}(M,TM)$ is a vector-valued $1$-form, $\ell\in\mathrm{End}(E)$ is an endomorphism of $E$, and $\nabla:\Gamma(E)\to\Omega^{1}(M,E)$ is an $\mathbb{R}$-linear map satisfying the Leibniz rule
$\nabla_{v}(f\xi)=f\nabla_{v}(\xi)+\left(\mathscr{L}_{v}f\right)\ell(\xi)-\left(\mathscr{L}_{r(v)}f\right)\xi,$ (18)
@@ 3.1 Characterization of IM tensor fields via 1-derivations
Correspondence between linear tensor fields and $1$-derivations.
@@
$\xi^{\uparrow}(e)\coloneqq\left.\frac{\mathrm{d}}{\mathrm{d}\tau}\right|_{\tau=0}\left(e+\tau\xi_{\pi(e)}\right).$ (19)
Linear $(1,1)$-tensor fields $R:TE\to TE$ are in one-to-one correspondence with $1$-derivations $\mathscr{D}=\left(\nabla,\ell,r\right)$ through the relations [9, Corollary 4.11]
$\nabla_{\mathrm{d}\pi(v)}(\xi)=\left(\mathscr{L}_{\xi^{\uparrow}}R\right)(v),\quad R\left(\xi^{\uparrow}\right)=\ell(\xi)^{\uparrow}\quad\text{and}\quad\left\langle R,\pi^{*}\alpha\right\rangle=\pi^{*}\left\langle r,\alpha\right\rangle$ (20)
@@
The Nijenhuis condition in terms of 1-derivations.
We can express the vanishing of the Nijenhuis torsion for a linear tensor field $R\in\Omega^{1}(E,TE)$ in terms of $1$-derivations. The tensor field $R$ is a Nijenhuis tensor field if and only if its corresponding $1$-derivation $(\nabla,\ell,r)$ satisfies the equations
$\begin{cases}\mathscr{N}_{r}(u,v)=0,\\ \ell\left(\nabla_{v}(\xi)\right)-\nabla_{v}\left(\ell(\xi)\right)=0,\\ \nabla^{2}_{(u,v)}(\xi)\coloneqq\ell\left(\nabla_{\left[u,v\right]}(\xi)\right)-\left[\nabla_{u},\nabla_{v}\right](\xi)-\nabla_{\left[u,v\right]_{r}}(\xi)=0,\end{cases}$ (21)
@@
Infinitesimally multiplicative 1-derivations.
The IM condition for $(1,1)$-tensor fields can be expressed equivalently in the language of $1$-derivations as follows. A $1$-derivation $\mathscr{D}=\left(\nabla,\ell,r\right)$ is said to be a IM $1$-derivation if it satisfies the so-called IM-equations [9, Section 6]:
$\begin{cases}\nabla_{x}\left(\left[\xi,\zeta\right]_{A}\right)=\left[\nabla_{v}(\xi),\zeta\right]_{A}+\left[\xi,\nabla_{v}(\zeta)\right]_{A}-\nabla_{\left[\rho(\zeta),v\right]_{A}}(\xi)-\nabla_{\left[\rho(\xi),v\right]_{A}}(\zeta)\\ \ell\left(\left[\xi,\zeta\right]_{A}\right)=\left[\xi,\ell(\zeta)\right]_{A}-\nabla_{\rho(\zeta)}(\xi)\\ \rho\left(\nabla_{v}(\xi)\right)=\nabla^{r}_{v}\left(\rho(\xi)\right),\\ r\circ\rho=\rho\circ\ell.\end{cases}$ (22)
An infinitesimal Nijenhuis structure on a Lie algebroid $A\to M$ can now be reformulated as an IM $1$-derivation $\mathscr{D}=(\nabla,\ell,r)$ satisfying the Nijenhuis equations (21).
@@
The tangent lift $r^{\mathrm{tg}}\in\Omega^{1}(TM,T(TM))$ of a given $(1,1)$-form $r\in\Omega^{1}(M,TM)$ is a linear $(1,1)$-tensor field on $TM\to M$ (Equation (25)) is always a IM tensor field ([16, Corollary 3.5]). We consider also the cotangent lift $r^{\mathrm{cotg}}\in\Omega^{1}(T^{*}M,T(T^{*}M))$, it is a linear tensor field constructed as follows. Let $\omega_{\mathrm{can}}\in\Omega^{2}(T^{*}M)$ be the canonical symplectic form on $T^{*}M$ and $\phi_{r}\coloneqq r^{*}:T^{*}M\to T^{*}M$. Then $r^{\mathrm{cotg}}$ is defined by
@@
If $r$ is a Nijenhuis tensor field, then both $r^{\mathrm{tg}}$ and $r^{\mathrm{cotg}}$ are Nijenhuis tensor fields. To define the corresponding $1$-derivations of $r^{\mathrm{tg}}$ and $r^{\mathrm{cotg}}$ consider the non-linear operators $\nabla^{r}:\mathfrak{X}(M)\to\Omega^{1}(M,TM)$ and $\nabla^{r,*}:\Omega^{1}(M)\to\Omega^{1}(M,T^{*}M)$ given by
@@
More explicitly, $\nabla^{r}_{v}(u)=\left[u,r(v)\right]-r\left(\left[u,v\right]\right)$ and $\nabla^{r,*}_{v}(\alpha)=\iota_{v}\mathrm{d}(r^{*}\alpha)-\iota_{r(v)}\mathrm{d}\alpha$. The corresponding $1$-derivations to the tangent and contangent lifts are
@@
respectively [16, Section 3].
@@
A holomorphic vector bundle can be characterized as a (real) vector bundle $E\to M$, together with a complex structure $r\in\Omega^{1}(M,TM)$ on $M$, an endomorphism $\ell:E\to E$ such that $\ell^{2}=-\mathrm{id}_{E}$ and a flat $T^{(1,0)}M$-connection $\nabla^{\prime}:\Gamma\left(T^{(1,0)}M\right)\times\Gamma(E)\to\Gamma(E)$ on the complex vector bundle $(E,\ell)$. Equivalently, the holomorphic structure of $E$ is encoded in the $1$-derivation $\mathscr{D}^{\mathrm{hol}}=\left(\nabla^{\mathrm{hol}},\ell,r\right)$, where
Moreover, a $1$-derivation $\mathscr{D}=\left(\nabla,\ell,r\right)$ defines a holomorphic vector bundle if and only if
$r^{2}=-\mathrm{id}_{TM},\quad\ell^{2}=-\mathrm{id}_{E}\quad\text{and}\quad\ell(\nabla_{v}(u))+\nabla_{r(v)}(u)=0,$ (23)
and the Nijenhuis conditions (21) are satisfied (see [8, Example 2.3]).
Example 3.4.
A holomorphic Lie algebroid [10, 38] is a holomorphic vector bundle $A\to M$ equipped with a holomorphic anchor and a Lie bracket that satisfies the Leibniz identity. Equivalently, it is a (real) Lie algebroid equipped with an IM $(1,1)$-tensor field $I_{A}\in\Omega^{1}(A,TA)$ with vanishing Nijenhuis torsion [25, Proposition 2.3]. In the language of $1$-derivations, the holomorphic structure $I_{A}$ corresponds to a IM $1$-derivation $(\nabla,\ell,r)$ satisfying (23).
3.2 The Lie correspondence for (1,1)-tensor fields
Let $\mathscr{G}$ be a Lie groupoid with Lie algebroid $A_{\mathscr{G}}$. The tangent Lie algebroid $TA_{\mathscr{G}}\to TM$ is integrable to the tangent Lie groupoid $T\mathscr{G}$ (specifically, via the canonical identification $TA_{\mathscr{G}}\cong A_{T\mathscr{G}}$; see [26, Theorem 7.1]). Assuming that $\mathscr{G}$ is source-simply connected, Lie’s second theorem guarantees that any IM $(1,1)$-tensor field $R:TA_{\mathscr{G}}\to TA_{\mathscr{G}}$, viewed as a Lie algebroid morphism, integrates to a unique Lie groupoid morphism, yielding a multiplicative $(1,1)$-tensor field $R_{\mathscr{G}}\in\Omega^{1}(\mathscr{G},T\mathscr{G})$. The Lie functor is the inverse of this integration process, establishing a bijection between multiplicative $(1,1)$-tensor fields and IM tensor fields. Furthermore, this bijection preserves the Nijenhuis condition on the $(1,1)$-tensor fields (see [25, Theorem 3.14]).

### v2

@@ 2.3. Morita equivalence of Nijenhuis groupoids
is a cartesian (pullback) diagram, i.e., $H_{1}\cong(\varphi\times\varphi)^{*}G_{1}$. Two Lie groupoids $\mathscr{G}$ and $\mathscr{H}$ are Morita equivalent if and only if they are connected by a span of Morita morphisms $\mathscr{G}\leftarrow\mathscr{P}\to\mathscr{H}$. The diagram $\mathscr{G}\leftarrow\mathscr{P}\to\mathscr{H}$ is also referred as a Morita equivalence (in terms of Morita morphisms) between $\mathscr{G}$ and $\mathscr{H}$.
@@
    (3) $\left(\mathscr{G},N_{1}\right)\xleftarrow{\quad\varphi_{1}\quad}\left(\mathscr{P},K\right)\xrightarrow{\quad\varphi_{2}\quad}\left(\mathscr{H},N_{2}\right),$
@@
    (4) $\mathrm{d}\varphi_{1}\circ K=N_{1}\circ\mathrm{d}\varphi_{1}\qquad\text{and}\qquad\mathrm{d}\varphi_{2}\circ K=N_{2}\circ\mathrm{d}\varphi_{2}.$
@@
    (5) $\left(G_{1},N_{1}\right)$$\left(P,J\right)$$\left(H_{1},N_{2}\right)$$\left(G_{0},r_{1}\right)$$\left(H_{0},r_{2}\right),$$\scriptstyle{\mu_{1}}$$\scriptstyle{\mu_{2}}$
@@
    (6) $N_{1}\times J\times N_{2}\left(TG_{1}\times_{TG_{0}}TP\times_{TH_{0}}TH_{1}\right)\subseteq TG_{1}\times_{TG_{0}}TP\times_{TH_{0}}TH_{1}$
@@
    (7) $J\circ\mathrm{d}\mathscr{A}=\mathrm{d}\mathscr{A}\circ\left(N_{1}\times J\times N_{2}\right),$
@@
(8) $g\cdot\left(g^{\prime},p\right)=\left(gg^{\prime},p\right)\qquad\text{and}\qquad\left(g,p^{\prime}\right)\cdot p=\left(g\,\varphi_{1}(p),\mathbf{s}_{\mathscr{P}}(p)\right).$
@@
(9) $\mathrm{d}\mathscr{A}_{1}\circ N_{1}\times(N_{1}\times r_{K})=\left(N_{1}\times r_{K}\right)\circ\mathrm{d}\mathscr{A}_{1}\qquad\text{and}$
(10) $\mathrm{d}\widetilde{\mathscr{A}}_{1}\circ\left(N_{1}\times r_{K}\right)\times K=\left(N_{1}\times r_{K}\right)\circ\mathrm{d}\widetilde{\mathscr{A}}_{1}.$
@@
(11) $p\cdot(p^{\prime},h)=\left(\mathbf{t}_{\mathscr{P}}(p),\varphi_{2}(p)\,h\right)\qquad\text{and}\qquad(p,h^{\prime})\cdot h=(p,h^{\prime}h).$
@@
(12) $\mathrm{d}\widetilde{\mathscr{A}}_{2}\circ\left(K\times\left(r_{K}\times N_{2}\right)\right)=\left(r_{K}\times N_{2}\right)\circ\mathrm{d}\widetilde{\mathscr{A}}_{2}\qquad\text{and}$
(13) $\mathrm{d}\mathscr{A}_{2}\circ\left(\left(r_{K}\times N_{2}\right)\times N_{2}\right)=\left(r_{K}\times N_{2}\right)\circ\mathrm{d}\mathscr{A}_{2}.$
@@
Conversely, assume there exists a Morita bibundle (5) satisfying (6) and (7). Let $\mathscr{P}=\left(P_{1}\rightrightarrows P\right)$ be the Lie groupoid where $P_{1}\coloneqq G_{1}\times_{(\mathbf{s}_{\mathscr{G}},\mu_{1})}P\times_{(\mu_{2},\mathbf{t}_{\mathscr{H}})}H_{1}$. The source and target maps are $(g,p,h)\mapsto p$ and $(g,p,h)\mapsto g\cdot p\cdot h$, respectively, while the multiplication is $((g,p,h),(g^{\prime},p^{\prime},h^{\prime}))\mapsto(gg^{\prime},p^{\prime},hh^{\prime})$. Then the diagram
@@
(14) $\mathrm{d}\mu_{1}\circ J=r_{1}\circ\mathrm{d}\mu_{1}\qquad\text{and}\qquad\mathrm{d}\mu_{2}\circ J=r_{2}\circ\mathrm{d}\mu_{2}.$
@@
(15) $N_{1}\times J\left(TG_{1}\times_{TG_{0}}TP\right)\subseteq TG_{1}\times_{TG_{0}}TP,\qquad J\times N_{2}\left(TP\times_{TH_{0}}TH_{1}\right)\subseteq TP\times_{TH_{0}}TH_{1},$
@@
(16) $J\circ\mathrm{d}\mathscr{A}_{1}=\mathrm{d}\mathscr{A}_{1}\circ\left(N_{1}\times J\right),\qquad J\circ\mathrm{d}\mathscr{A}_{2}=\mathrm{d}\mathscr{A}_{2}\circ\left(J\times N_{2}\right).$
@@
2.4. Morita equivalence is an equivalence relation.
@@
(17) $P\diamond Q\coloneqq\frac{P\times_{H_{0}}Q}{\mathscr{H}},$
@@
For reflexivity, let $\mathscr{G}$ be a Lie groupoid equipped with a multiplicative $(1,1)$-tensor field $N\in\Omega^{1}(\mathscr{G},T\mathscr{G})$. Because the actions of the principal bibundle $G_{0}\xleftarrow{\ \mathbf{t}\ }G_{1}\xrightarrow{\ \mathbf{s}\ }G_{0}$ correspond to left and right multiplication, and because NN is multiplicative, the tensor fields NN and $N\times N\times N$ satisfy (7).
@@
3. Infinitesimal Nijenhuis structures
In this section, we study the infinitesimal version of the notion of Morita equivalence for Nijenhuis groupoids that we developed in Section 2. We begin by defining infinitesimal Nijenhuis structures and providing an equivalent characterization in terms of 11-derivations. We define the notion of Morita equivalence of infinitesimal Nijenhuis structures in Definition 3.13, and in Theorem 3.19 we show that it corresponds to the differentiation of the global Morita equivalence of Nijenhuis groupoids.
@@ 3. Infinitesimal Nijenhuis structures
An infinitesimal Nijenhuis structure is a pair $(A,R)$, where $A\to M$ is a Lie algebroid and $R\in\Omega^{1}(A,TA)$ is a Nijenhuis tensor field on the total space of AA, such that $R:TA\to TA$ is a morphism of Lie algebroids from the tangent Lie algebroid to itself.
@@
for some $r\in\Omega^{1}(M,TM)$. Linear $(1,1)$-tensor fields that are morphisms of Lie algebroids are the infinitesimal counterparts of multiplicative $(1,1)$-tensor fields, they are called IM (infinitesimally multiplicative) $(1,1)$-tensor fields [undefg]. These objects can be characterized in algebraic terms as 11-derivations as we explain in the following subsection.
3.1. Characterization of IM tensor fields via 1-derivations
A 11-derivation on a vector bundle $\pi:E\to M$ is a triple $\mathscr{D}=\left(\nabla,\ell,r\right)$, where $r\in\Omega^{1}(M,TM)$ is a vector-valued 11-form, $\ell\in\mathrm{End}(E)$ is an endomorphism of $E$, and $\nabla:\Gamma(E)\to\Omega^{1}(M,E)$ is an $\mathbb{R}$-linear map satisfying the Leibniz rule
(18) $\nabla_{v}(f\xi)=f\nabla_{v}(\xi)+\left(\mathscr{L}_{v}f\right)\ell(\xi)-\left(\mathscr{L}_{r(v)}f\right)\xi,$
@@ 3.1. Characterization of IM tensor fields via 1-derivations
3.1.1. Correspondence between linear tensor fields and 11-derivations
@@
(19) $\xi^{\uparrow}(e)\coloneqq\left.\frac{\mathrm{d}}{\mathrm{d}\tau}\right|_{\tau=0}\left(e+\tau\xi_{\pi(e)}\right).$
Linear $(1,1)$-tensor fields $R:TE\to TE$ are in one-to-one correspondence with 11-derivations $\mathscr{D}=\left(\nabla,\ell,r\right)$ through the relations [undefg, Corollary 4.11]
(20) $\nabla_{\mathrm{d}\pi(v)}(\xi)=\left(\mathscr{L}_{\xi^{\uparrow}}R\right)(v),\quad R\left(\xi^{\uparrow}\right)=\ell(\xi)^{\uparrow}\quad\text{and}\quad\left\langle R,\pi^{*}\alpha\right\rangle=\pi^{*}\left\langle r,\alpha\right\rangle,$
@@
3.1.2. The Nijenhuis condition in terms of 1-derivations
We can express the vanishing of the Nijenhuis torsion for a linear tensor field $R\in\Omega^{1}(E,TE)$ in terms of 11-derivations. The tensor field $R$ is a Nijenhuis tensor field if and only if its corresponding 11-derivation $(\nabla,\ell,r)$ satisfies the equations
(21) $\begin{cases}\mathscr{N}_{r}(u,v)=0,\\ \ell\left(\nabla_{v}(\xi)\right)-\nabla_{v}\left(\ell(\xi)\right)=0,\\ \nabla^{2}_{(u,v)}(\xi)\coloneqq\ell\left(\nabla_{\left[u,v\right]}(\xi)\right)-\left[\nabla_{u},\nabla_{v}\right](\xi)-\nabla_{\left[u,v\right]_{r}}(\xi)=0,\end{cases}$
@@ 3.1.2. The Nijenhuis condition in terms of 1-derivations
3.1.3. Infinitesimally multiplicative 1-derivations
The IM condition for $(1,1)$-tensor fields can be expressed equivalently in the language of 11-derivations as follows. A 11-derivation $\mathscr{D}=\left(\nabla,\ell,r\right)$ is said to be an IM 11-derivation if it satisfies the so-called IM-equations [undefg, Section 6]:
(22) $\begin{cases}\nabla_{v}\left(\left[\xi,\zeta\right]_{A}\right)=\left[\nabla_{v}(\xi),\zeta\right]_{A}+\left[\xi,\nabla_{v}(\zeta)\right]_{A}-\nabla_{\left[\rho(\zeta),v\right]_{A}}(\xi)-\nabla_{\left[\rho(\xi),v\right]_{A}}(\zeta)\\ \ell\left(\left[\xi,\zeta\right]_{A}\right)=\left[\xi,\ell(\zeta)\right]_{A}-\nabla_{\rho(\zeta)}(\xi)\\ \rho\left(\nabla_{v}(\xi)\right)=\nabla^{r}_{v}\left(\rho(\xi)\right),\\ r\circ\rho=\rho\circ\ell.\end{cases}$
An infinitesimal Nijenhuis structure on a Lie algebroid $A\to M$ can now be reformulated as an IM 11-derivation $\mathscr{D}=(\nabla,\ell,r)$ satisfying the Nijenhuis equations (21).
@@ 3.1.3. Infinitesimally multiplicative 1-derivations
The tangent lift $r^{\mathrm{tg}}\in\Omega^{1}(TM,T(TM))$ of a given $(1,1)$-form $r\in\Omega^{1}(M,TM)$ is a linear $(1,1)$-tensor field on $TM\to M$ (Equation (25)) is always an IM tensor field ([undefp, Corollary 3.5]). We consider also the cotangent lift $r^{\mathrm{cotg}}\in\Omega^{1}(T^{*}M,T(T^{*}M))$, it is a linear tensor field constructed as follows. Let $\omega_{\mathrm{can}}\in\Omega^{2}(T^{*}M)$ be the canonical symplectic form on $T^{*}M$ and $\phi_{r}\coloneqq r^{*}:T^{*}M\to T^{*}M$. Then $r^{\mathrm{cotg}}$ is defined by
@@
If $r$ is a Nijenhuis tensor field, then both $r^{\mathrm{tg}}$ and $r^{\mathrm{cotg}}$ are Nijenhuis tensor fields. To define the corresponding 11-derivations of $r^{\mathrm{tg}}$ and $r^{\mathrm{cotg}}$ consider the non-linear operators $\nabla^{r}:\mathfrak{X}(M)\to\Omega^{1}(M,TM)$ and $\nabla^{r,*}:\Omega^{1}(M)\to\Omega^{1}(M,T^{*}M)$ given by
@@
More explicitly, $\nabla^{r}_{v}(u)=\left[u,r(v)\right]-r\left(\left[u,v\right]\right)$ and $\nabla^{r,*}_{v}(\alpha)=\iota_{v}\mathrm{d}(r^{*}\alpha)-\iota_{r(v)}\mathrm{d}\alpha$. The corresponding 11-derivations to the tangent and cotangent lifts are
@@
respectively [undefp, Section 3].
Example $3.$
A holomorphic vector bundle can be characterized as a (real) vector bundle $E\to M$, together with a complex structure $r\in\Omega^{1}(M,TM)$ on $M$, an endomorphism $\ell:E\to E$ such that $\ell^{2}=-\mathrm{id}_{E}$ and a flat $T^{(1,0)}M$-connection $\nabla^{\prime}:\Gamma\left(T^{(1,0)}M\right)\times\Gamma(E)\to\Gamma(E)$ on the complex vector bundle $(E,\ell)$. Equivalently, the holomorphic structure of $E$ is encoded in the 11-derivation $\mathscr{D}^{\mathrm{hol}}=\left(\nabla^{\mathrm{hol}},\ell,r\right)$, where
Moreover, a 11-derivation $\mathscr{D}=\left(\nabla,\ell,r\right)$ defines a holomorphic vector bundle if and only if
(23) $r^{2}=-\mathrm{id}_{TM},\quad\ell^{2}=-\mathrm{id}_{E}\quad\text{and}\quad\ell(\nabla_{v}(u))+\nabla_{r(v)}(u)=0,$
and the Nijenhuis conditions (21) are satisfied (see [undefh, Example 2.3]).
@@
A holomorphic Lie algebroid [undefj, undefal] is a holomorphic vector bundle $A\to M$ equipped with a holomorphic anchor and a Lie bracket that satisfies the Leibniz identity. Equivalently, it is a (real) Lie algebroid equipped with an IM $(1,1)$-tensor field $I_{A}\in\Omega^{1}(A,TA)$ with vanishing Nijenhuis torsion [undefy, Proposition 2.3]. In the language of 11-derivations, the holomorphic structure $I_{A}$ corresponds to an IM 11-derivation $(\nabla,\ell,r)$ satisfying (23).
3.2. The Lie correspondence for (1,1)-tensor fields
Let $\mathscr{G}$ be a Lie groupoid with Lie algebroid $A_{\mathscr{G}}$. The tangent Lie algebroid $TA_{\mathscr{G}}\to TM$ is integrable to the tangent Lie groupoid $T\mathscr{G}$ (specifically, via the canonical identification $TA_{\mathscr{G}}\cong A_{T\mathscr{G}}$; see [undefaa, Theorem 7.1]). Assuming that $\mathscr{G}$ is source-simply connected, Lie’s second theorem guarantees that any IM $(1,1)$-tensor field $R:TA_{\mathscr{G}}\to TA_{\mathscr{G}}$, viewed as a Lie algebroid morphism, integrates to a unique Lie groupoid morphism, yielding a multiplicative $(1,1)$-tensor field $R_{\mathscr{G}}\in\Omega^{1}(\mathscr{G},T\mathscr{G})$. The Lie functor is the inverse of this integration process, establishing a bijection between multiplicative $(1,1)$-tensor fields and IM tensor fields. Furthermore, this bijection preserves the Nijenhuis condition on the $(1,1)$-tensor fields (see [undefy, Theorem 3.14]).